OpenStudy (anonymous):

Find the series of equations which are perpendicular to y=2x^2?

OpenStudy (anonymous):

I found the slope -1/4x

OpenStudy (tkhunny):

Excellent. Now, you need an antiderivative of that.

OpenStudy (anonymous):

That just means integrate it, right?

OpenStudy (anonymous):

@tkhunny

OpenStudy (tkhunny):

Mostly. Let's see what you get.

OpenStudy (anonymous):

\[-\frac{ x^{2} }{ 8 } +c\]

OpenStudy (tkhunny):

?? Let's clarify. You had \(y = 2x^{2}\). This gives \(\dfrac{dy}{dx} = 4x\). Therefore, the curevs we seek have \(\dfrac{dy}{dx} = -\dfrac{1}{4x}\), resulting in what antiderivative?

OpenStudy (anonymous):

I think I'm a little confused... I figured the antiderivative of -1/4x would be what I said

OpenStudy (tkhunny):

Better brush up on your exponents. Negative One (-1) is special! \(\int -\dfrac{1}{4x}\;dx = -\dfrac{1}{4}\int \dfrac{1}{x}\;dx = -\dfrac{1}{4}\cdot\ln(|x|) + C\)

OpenStudy (anonymous):

I do that every time! Would that be my equation? In my notes they have a y added and I'm not sure why